# Density-Aware Farthest Point Sampling

### v1

@@ Abstract
We focus on training machine learning regression models in scenarios where the availability of labeled training data is limited due to computational constraints or high labeling costs. Thus, selecting suitable training sets from unlabeled data is essential for balancing performance and efficiency. For the selection of the training data, we focus on passive and model-agnostic sampling methods that only consider the data feature representations. We derive an upper bound for the expected prediction error of Lipschitz continuous regression models that linearly depends on the weighted fill distance of the training set—a quantity we can estimate simply by considering the data features. We introduce "Density-Aware Farthest Point Sampling" (DA-FPS), a novel sampling method. We prove that DA-FPS provides approximate minimizers for a data-driven estimation of the weighted fill distance, thereby aiming at minimizing our derived bound. We conduct experiments using two regression models across three datasets. The results demonstrate that DA-FPS significantly reduces the mean absolute prediction error compared to other sampling strategies.
@@ 1 Introduction
Machine learning regression aims to predict continuous numerical values based on input features by learning the underlying prediction patterns from labeled training data. Data quality and data selection are crucial for developing and deploying effective regression models. Without sufficient or reliable data, ML regression models cannot perform well: too little data leads to poor learning, and biased or corrupted data results in inaccurate predictions. This work addresses challenges related to the limited availability of labeled training data, often resulting from computational constraints and high labeling costs. These challenges frequently occur in scientific applications, where labeling typically relies on expensive numerical simulations or laboratory experiments. Our work is motivated by molecular property prediction. The chemical compound space in large. It is estimated to contain over $10^{60}$ molecules (Kirkpatrick & Ellis, 2004). Its exhaustive exploration using classical labeling methods based on quantum-mechanics simulation is impractical. Machine learning regression offers a promising solution by leveraging small, selectively labeled datasets to predict molecular properties at scale (von Lilienfeld et al., 2020). This data-efficient approach has the potential to accelerate the discovery of new drugs and materials.
@@
We consider scenarios where we have access to a large pool of unlabeled data, e.g., molecules for which chemical and physical properties (the labels) are unknown and costly to obtain. We aim to select and label a small subset of the data pool to train a regression model. We show that by selecting suitable training sets we can positively impact the prediction performance of ML regression models. The goal is to improve the average prediction performance of the model on the points in the pool not selected for training. We focus on label-agnostic, passive, and model-agnostic sampling, that is, selection approaches that solely rely on the data feature representations, do not consider any active learning procedure, and do not assume any specific structure for the regression model. On the one hand, this promotes the reusability of the labeled samples, ensuring labeling efforts are not wasted on subsets useful only for specific models or tasks. On the other hand, developing such approaches is a non-trivial challenge because we can rely only on very little information, that is, the data feature representations and some assumption on the regularity of the learned function, but not on the model used to learn it, e.g., using a neural network or a kernel approach. As an additional constraint, we are interested on sampling approaches that are computationally efficient and potentially can be used on data pools consisting of hundreds of thousands of high-dimensional data points, which is a realistic size for molecular property prediction tasks.
@@
Training data selection for model performance optimization has been extensively studied. Relevant literature includes Active Learning (AL) (Settles, 2012; Ren et al., 2021) and experimental design (John & Draper, 1975; Yu et al., 2006). However, both focus on methodologies that optimize data selection by leveraging model-specific knowledge. Here, we focus on passive and model-agnostic sampling strategies. Passive sampling was first introduced in Yu & Kim (2010) where the authors highlight its importance in scenarios where labeled data can be difficult, time-consuming, or expensive to obtain. Drawing from Climaco & Garcke (2024), passive sampling strategies can be classified as model-dependent and model-agnostic. Model-dependent approaches optimize the data selection process for specific models or classes of models. We see single-shot batch active learners and experimental design approaches as passive addition model-dependent methods. Model-agnostic selection approaches have the potential to benefit multiple classes of regression models, thereby enhancing reusability of costly data labelling.
@@
The Farthest Point Sampling (FPS) is a greedy passive and model-agnostic sampling strategy that provides approximate solutions to the $k$-center problem by minimizing the fill distance of the selected set, which is the maximal distance between a point in the feature space of interest and its closest selected point. The benefits of minimizing the training set fill distance have been studied independently for classification (Sener & Savarese, 2018) and regression (Climaco & Garcke, 2024). In Sener & Savarese (2018) it was shown that minimizing the training set fill distance leads to increased accuracy by reducing the average classification error of Lipschitz continuous models with a softmax output layer. Climaco & Garcke (2024) show that for regression tasks minimizing the training set fill distance with the FPS significantly reduces the maximum error of the label predictions of Lipschitz continuous models. However, it was also shown that employing FPS to minimize the maximum prediction error is mostly beneficial in the low data regime and that in such low data regime it does not provide any significant advantage in terms of the average absolute prediction error compared to other passive sampling approaches that are model-agnostic.
@@
This work addresses the limitations of the FPS for molecular property prediction tasks and aims to extend the work done in Climaco & Garcke (2024). In particular, we develop a passive and model-agnostic algorithm, inspired by the FPS greedy procedure and supported by a solid theoretical motivation, to select training sets that can improve the average prediction performances of regression models measured in terms of the mean absolute error (MAE).
@@
We derive an upper bound for the expected prediction error of Lipschitz continuous regression models. Our bound is linear in a novel weighted fill distance, that is, the maximum weighted distance between a point in the feature domain of interest and its closest selected training point. The weights we use to scale the distances provide information on the distributions of the data in the feature space. We introduce a novel passive and model-agnostic data selection strategy called “Density-Aware Farthest Point Sampling” (DA-FPS). This strategy greedily selects sets from a pool of available data points with the aim of minimizing a data-drive estimation of the weighted fill distance of the selected set, thereby minimizing our proposed bound. The ultimate goal of DA-FPS is to select training sets that can enhance the average prediction performance of Lipschitz continuous regression models. We prove that DA-FPS provides approximate minimizers for a data-driven estimation of the weighted fill distance. Additionally, we report experimental results showing that selecting training sets using DA-FPS can decrease the average prediction absolute error of Lipschitz continuous regression models. Compared with other model-agnostic sampling techniques, including FPS, DA-FPS demonstrates its superiority for low training set budgets in terms of mean absolute error (MAE) reduction. We are not aware of any other passive sampling model-agnostic strategy supported by rigorous theoretical work that can improve the mean absolute prediction error of a regression model.
@@ 2 Related work
There has been a significant research effort to develop approaches that can be used for training data selection aiming at model performance optimization. There is a large body of literature on active learning (AL). See Settles (2012) and Ren et al. (2021) for a review of such approaches. AL typically involves training regression models to predict uncertainties or estimate labels for unlabeled data, selecting the most relevant points for labeling, including them in the training set and repeating the process until stopping criteria are met. It optimizes the training set selection for a specific model or model class, and for a given task, by using computed labels to iteratively update model parameters during selection. Single-shot batch active learners select the training set in one step without iterative training or label knowledge. Examples include methods that minimize maximum mean discrepancy, discrepancy, and nuclear discrepancy, which are quantities that can be estimated solely relying on the data features (Chattopadhyay et al., 2012; Wang & Ye, 2015; Viering et al., 2019). However, these approaches require choosing a kernel function, and their effectiveness in reducing the prediction error is theoretically motivated only assuming that the regression function belongs to the reproducing kernel Hilbert space associated with the chosen kernel (Viering et al., 2019). That is, they optimize the selection for specific model classes, such as kernel regularized least squares, similar to active learning. Moreover, such approaches have been mostly studied for classification tasks. Assuming the knowledge of the learning model may even lead to the development of strategies that reflect some principle of optimality for the selection of the training set, as in the case of approaches developed in the field of experimental design (John & Draper, 1975; Yu et al., 2006).
@@
In this work we focus on passive and model-agnostic sampling strategies. Coresets (Feldman, 2019) provide an effective set of tools for performing model-agnostic passive sampling. Uniform random sampling is considered the natural benchmark for all the other sampling strategies (Feldman, 2019). Importance sampling selects data points that are more important to the problem at hand based on importance weights determining the relevance of each data point. For instance, Xie et al. (2023) present an importance sampling approach for selecting suitable pretraining datasets for large language models. Cluster-based methods, such as $k$-medoids++ and $k$-means++ (Schubert & Rousseeuw, 2021), group similar points together and select representatives from each cluster. Greedy methods such as submodular function maximization algorithms (Krause & Golovin, 2014) start with an empty set and iteratively add the most influential point at each step, according to a given principle. Data twinning partitions the dataset into statistically similar twin set by attempting to minimizing the energy distance, a quantity that can be estimated solely computing pairwise distances between data features. It has been effectively used for training data selection in regression (Vakayil & Joseph, 2022). The work most similar to ours is Climaco & Garcke (2024), where the authors show that selecting the training set by fill distance minimization using the Farthest Point Sampling (FPS) can reduce the maximum prediction error of Lipschitz continuous regression models. However, the authors also show that FPS does not offer any significant benefit in terms of the average absolute prediction error when compared to other benchmark passive sampling model-agnostic approaches.
@@ 3 Problem definition
Our setting is similar to the one introduced in Climaco & Garcke (2024). We consider a supervised regression problem where the feature space is $\mathcal{X}\subset\mathbb{R}^{d}$, bounded, and the label space is $\mathcal{Y}\subset\mathbb{R}$. The goal is to learn a function $g:\mathcal{X}\rightarrow\mathcal{Y}$ from a function space $\mathcal{M}$, using a set of labeled data. Each function in $\mathcal{M}$ may be parameterized by weights ${\bm{w}}\in\mathbb{R}^{m}$, depending on the learning algorithm. To evaluate prediction quality, we use an error function $l:\mathcal{X}\times\mathcal{Y}\times\mathcal{M}\rightarrow\mathbb{R}^{+}$, which measures the prediction error of a trained regression model for a given input feature and associated true label.
@@
Let $\mathcal{D}:=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ be a pool of available data points consisting of i.i.d. realizations of two random variables $\bm{X}$} and $Y$ taking value on $\mathcal{X}$ and $\mathcal{Y}$, respectively, with joint distribution $p_{\mathcal{D}}\in\mathcal{P}:=\{p:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}|\;\int_{\mathcal{X}\times\mathcal{Y}}p({\bm{x}},y)d{\bm{x}}dy=1\}$. We consider scenarios where the labels $\{y_{i}\}_{i=1}^{n}$ are unknown. We aim to select and label a training set $\mathcal{L}:=\{({\bm{x}}_{i_{j}},y_{i_{j}})\}_{j=1}^{b}$ from $\mathcal{D}$, with $i_{j}\in\{1,\dots,n\}\;\forall j$, and $b\ll n$. The objective of the training data selection is to maximize the average predictive performance of a regression model $m_{\mathcal{L}}\in\mathcal{M}$, trained on $\mathcal{L}$, on the new data points not selected for training. In the following, without loss of generality, we assume $\mathcal{L}:=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{b}$ and define $\mathcal{L}_{\mathcal{X}}:=\{{\bm{x}}_{j}\}_{j=1}^{b}$. Following along Wang & Ye (2015), we assume that the training data is selected according to a distribution $p_{\mathcal{L}}\in\mathcal{P}$ with $p_{\mathcal{D}}\neq p_{\mathcal{L}}$ and $p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. In other words, the training data distribution may differ from the data source distribution, but the map connecting a data location ${\bm{x}}\in\mathcal{X}$ and its associated label value is independent on how the data is selected. Furthermore, we study a typical scenario involving molecular prediction tasks in which the data source distribution $p_{\mathcal{D}}$ is unknown, and we have only access to the locations $\mathcal{D}_{\mathcal{X}}:=\{{\bm{x}}_{i}\}_{i=1}^{n}$ of the dataset $\mathcal{D}$. Moreover, we are interested in applications in which the labeling process is computationally expensive, therefore, given a budget $b\in\mathbb{N}$ of points to label, the goal is to select a subset $\mathcal{L}_{\mathcal{X}}\subset\mathcal{D}_{\mathcal{X}}$ and label it to obtain a subset $\mathcal{L}\subset\mathcal{D}$ that solves the following optimization problem:
$\min_{\begin{subarray}{c}\mathcal{L}\subset\mathcal{D},\\ |\mathcal{L}|=b\end{subarray}}\mathbb{E}_{p_{\mathcal{D}}}[l({\bm{X}},Y,m_{\mathcal{L}})],$ (1)
@@
The optimization problem addressed in Climaco & Garcke (2024) and the one we define in (1) are similar in the sense that both focus on minimizing an expected value of the error function by selecting a subset $\mathcal{L}$ from a set of available data points $\mathcal{D}$. The key distinction lies in their objectives: in  Climaco & Garcke (2024), the goal is to minimize the maximum expected error over a finite set of known data locations. Differently, in (1), the objective is to minimize the average error, defined as the expected value of the error function over the data source distribution $p_{\mathcal{D}}$ on $\mathcal{X}\times\mathcal{Y}$.

### v2

@@ Abstract
We focus on training machine learning regression models in scenarios where the availability of labeled training data is limited due to computational constraints or high labeling costs. Thus, selecting suitable training sets from unlabeled data is essential for balancing performance and efficiency. For the selection of the training data, we focus on passive and model-agnostic sampling methods that only consider the data feature representations. We derive an upper bound for the expected prediction error of Lipschitz continuous regression models that linearly depends on the weighted fill distance of the training set—a quantity we can estimate simply by considering the data features. We introduce “Density-Aware Farthest Point Sampling” (DA-FPS), a novel sampling method. We prove that DA-FPS provides approximate minimizers for a data-driven estimation of the weighted fill distance, thereby aiming at minimizing our derived bound. We conduct experiments using two regression models across three datasets. The results demonstrate that DA-FPS significantly reduces the mean absolute prediction error compared to other sampling strategies.
@@ 1 Introduction
Machine learning regression aims to predict continuous numerical values based on input features by learning the underlying prediction patterns from labeled training data. Data quality and data selection are crucial for developing and deploying effective regression models. Without sufficient or reliable data, ML regression models cannot perform well: too little data leads to poor learning, and biased or corrupted data results in inaccurate predictions. This work addresses challenges related to the limited availability of labeled training data, often resulting from computational constraints and high labeling costs. These challenges frequently occur in scientific applications, where labeling typically relies on expensive numerical simulations or laboratory experiments. Our work is motivated by molecular property prediction. The chemical compound space is large. It is estimated to contain over $10^{60}$ molecules (Kirkpatrick and Ellis, 2004). Its exhaustive exploration using classical labeling methods based on quantum-mechanics simulation is impractical. Machine learning regression offers a promising solution by leveraging small, selectively labeled datasets to predict molecular properties at scale (von Lilienfeld et al., 2020). This data-efficient approach has the potential to accelerate the discovery of new drugs and materials.
@@
We consider scenarios where we have access to a large pool of unlabeled data, e.g., molecules for which chemical and physical properties (the labels) are unknown and costly to obtain. We aim to select and label a small subset of the data pool to train a regression model. We show that by selecting suitable training sets we can positively impact the prediction performance of ML regression models. The goal is to improve the average prediction performance of the model on the points in the pool not selected for training. We focus on label-agnostic, passive, and model-agnostic sampling, that is, selection approaches that solely rely on the data feature representations, do not consider any active learning procedure, and do not assume any specific structure for the regression model. On the one hand, this promotes the reusability of the labeled samples, ensuring labeling efforts are not wasted on subsets useful only for specific models or tasks. On the other hand, developing such approaches is a non-trivial challenge because we can rely only on very little information, that is, the data feature representations and some assumption on the regularity of the learned function, but not on the model used to learn it, e.g., using a neural network or a kernel approach. As an additional constraint, we are interested in sampling approaches that are computationally efficient and potentially can be used on data pools consisting of hundreds of thousands of high-dimensional data points, which is a realistic size for molecular property prediction tasks.
@@
Training data selection for model performance optimization has been extensively studied. Relevant literature includes Active Learning (AL) (Settles, 2012; Ren et al., 2021) and experimental design (John and Draper, 1975; Yu et al., 2006). However, both focus on methodologies that optimize data selection by leveraging model-specific knowledge. Here, we focus on passive and model-agnostic sampling strategies. Passive sampling was first introduced in Yu and Kim (2010), where the authors highlight its importance in scenarios where labeled data can be difficult, time-consuming, or expensive to obtain. Drawing from Climaco and Garcke (2024), passive sampling strategies can be classified as model-dependent and model-agnostic. Model-dependent approaches optimize the data selection process for specific models or classes of models. We see single-shot batch active learners and experimental design approaches as passive and model-dependent methods. Model-agnostic selection approaches have the potential to benefit multiple classes of regression models, thereby enhancing reusability of costly data labelling.
@@
The Farthest Point Sampling (FPS) is a greedy passive and model-agnostic sampling strategy that provides approximate solutions to the $k$-center problem by minimizing the fill distance of the selected set, which is the maximal distance between a point in the feature space of interest and its closest selected point. The benefits of minimizing the training set fill distance have been studied independently for classification (Sener and Savarese, 2018) and regression (Climaco and Garcke, 2024). In Sener and Savarese (2018) it was shown that minimizing the training set fill distance leads to increased accuracy by reducing the average classification error of Lipschitz continuous models with a softmax output layer. In Climaco and Garcke (2024) we show that for regression tasks minimizing the training set fill distance with the FPS significantly reduces the maximum error of the label predictions of Lipschitz continuous models. However, it was also shown that employing FPS to minimize the maximum prediction error is mostly beneficial in the low data regime and that in such low data regime it does not provide any significant advantage in terms of the average absolute prediction error compared to other passive sampling approaches that are model-agnostic.
@@
This work addresses the limitations of the FPS for molecular property prediction tasks and aims to extend the work done in Climaco and Garcke (2024). In particular, we develop a passive and model-agnostic algorithm, inspired by the FPS greedy procedure and supported by a solid theoretical motivation, to select training sets that can improve the average prediction performance of regression models measured in terms of the mean absolute error (MAE).
@@
We derive an upper bound for the expected prediction error of Lipschitz continuous regression models. Our bound is linear in a novel weighted fill distance, that is, the maximum weighted distance between a point in the feature domain of interest and its closest selected training point. The weights we use to scale the distances provide information on the distributions of the data in the feature space. We introduce a novel passive and model-agnostic data selection strategy called “Density-Aware Farthest Point Sampling” (DA-FPS). This strategy greedily selects sets from a pool of available data points with the aim of minimizing a data-driven estimation of the weighted fill distance of the selected set, thereby minimizing our proposed bound. The ultimate goal of DA-FPS is to select training sets that can enhance the average prediction performance of Lipschitz continuous regression models. We prove that DA-FPS provides approximate minimizers for a data-driven estimation of the weighted fill distance. Additionally, we report experimental results showing that selecting training sets using DA-FPS can decrease the average prediction absolute error of Lipschitz continuous regression models. Compared with other model-agnostic sampling techniques, including FPS, DA-FPS demonstrates its superiority for low training set budgets in terms of mean absolute error (MAE) reduction. We are not aware of any other passive sampling model-agnostic strategy supported by rigorous theoretical work that can improve the mean absolute prediction error of a regression model.
The paper is organized as follows: In section 2 we review related work. In section 3 we formalize the regression problem and define the optimization objective. In section 4 we derive an upper bound for the expected prediction error of Lipschitz continuous regression models that depends linearly on a weighted fill distance of the training set. In section 5-7 we introduce the DA-FPS algorithm and prove that it provides approximate minimizers for a data-driven estimation of the weighted fill distance. In section 8 we report experimental results showing that DA-FPS can reduce the mean absolute prediction error of regression models compared to other passive sampling model-agnostic strategies. In section 9 we summarize our contributions and outline future research directions. After the conclusion, Table 1 provides key notation.
@@ 2 Related work
There has been a significant research effort to develop approaches that can be used for training data selection aiming at model performance optimization. There is a large body of literature on active learning (AL). See Settles (2012) and Ren et al. (2021) for a review of such approaches. AL typically involves training regression models to predict uncertainties or estimate labels for unlabeled data, selecting the most relevant points for labeling, including them in the training set and repeating the process until stopping criteria are met. It optimizes the training set selection for a specific model or model class, and for a given task, by using computed labels to iteratively update model parameters during selection. Single-shot batch active learners select the training set in one step without iterative training or label knowledge. Examples include methods that minimize maximum mean discrepancy, discrepancy, and nuclear discrepancy, which are quantities that can be estimated solely relying on the data features (Chattopadhyay et al., 2012; Wang and Ye, 2015; Viering et al., 2019). However, these approaches require choosing a kernel function, and their effectiveness in reducing the prediction error is theoretically motivated only assuming that the regression function belongs to the reproducing kernel Hilbert space associated with the chosen kernel (Viering et al., 2019). That is, they optimize the selection for specific model classes, such as kernel regularized least squares, similar to active learning. Moreover, such approaches have been mostly studied for classification tasks. Assuming the knowledge of the learning model may even lead to the development of strategies that reflect some principle of optimality for the selection of the training set, as in the case of approaches developed in the field of experimental design (John and Draper, 1975; Yu et al., 2006).
@@
In this work we focus on passive and model-agnostic sampling strategies. Coresets (Feldman, 2019) provide an effective set of tools for performing model-agnostic passive sampling. Uniform random sampling is considered the natural benchmark for all the other sampling strategies (Feldman, 2019). Importance sampling selects data points that are more important to the problem at hand based on importance weights determining the relevance of each data point. For instance, Xie et al. (2023) present an importance sampling approach for selecting suitable pretraining datasets for large language models. Cluster-based methods, such as $k$-medoids++ and $k$-means++ (Schubert and Rousseeuw, 2021), group similar points together and select representatives from each cluster. Greedy methods such as submodular function maximization algorithms (Krause and Golovin, 2014) start with an empty set and iteratively add the most influential point at each step, according to a given principle. Data twinning partitions the dataset into statistically similar twin sets by attempting to minimize the energy distance, a quantity that can be estimated solely by computing pairwise distances between data features. It has been effectively used for training data selection in regression (Vakayil and Joseph, 2022). In our earlier work Climaco and Garcke (2024) we show that selecting the training set by fill distance minimization using the Farthest Point Sampling (FPS) can reduce the maximum prediction error of Lipschitz continuous regression models. However, we also show that FPS does not offer any significant benefit in terms of the average absolute prediction error when compared to other benchmark passive sampling model-agnostic approaches.
@@ 3 Problem definition
In this section we formalize the setting of the regression problem we aim to address and define the optimization objective. Our setting is similar to the one introduced in Climaco and Garcke (2024). We consider a supervised regression problem where the feature space $\mathcal{X}\subset\mathbb{R}^{d}$ is bounded and the label space is $\mathcal{Y}\subset\mathbb{R}$. The goal is to learn a function $g:\mathcal{X}\rightarrow\mathcal{Y}$ from a function space $\mathcal{M}$, using a set of labeled data. Each function in $\mathcal{M}$ may be parameterized by weights ${\bm{w}}\in\mathbb{R}^{m}$, depending on the learning algorithm. To evaluate prediction quality, we use an error function $l:\mathcal{X}\times\mathcal{Y}\times\mathcal{M}\rightarrow\mathbb{R}^{+}$, which measures the prediction error of a trained regression model for a given input feature and associated true label.
@@
Let $\mathcal{D}:=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ be a pool of available data points consisting of i.i.d. realizations of two random variables $\bm{X}$} and $Y$ taking value on $\mathcal{X}$ and $\mathcal{Y}$, respectively, with joint distribution $p_{\mathcal{D}}\in\mathcal{P}:=\{p:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}|\;\int_{\mathcal{X}\times\mathcal{Y}}p({\bm{x}},y)d{\bm{x}}dy=1\}$. We consider scenarios where the labels $\{y_{i}\}_{i=1}^{n}$ are unknown. We aim to select and label a training set $\mathcal{L}:=\{({\bm{x}}_{i_{j}},y_{i_{j}})\}_{j=1}^{b}$ from $\mathcal{D}$, with $i_{j}\in\{1,\dots,n\}\;\forall j$, and $b\ll n$. The objective of the training data selection is to maximize the average predictive performance of a regression model $m_{\mathcal{L}}\in\mathcal{M}$, trained on $\mathcal{L}$, on the new data points not selected for training. In the following, without loss of generality, we assume $\mathcal{L}:=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{b}$ and define $\mathcal{L}_{\mathcal{X}}:=\{{\bm{x}}_{j}\}_{j=1}^{b}$. Following along Wang and Ye (2015), we assume that the training data is selected according to a distribution $p_{\mathcal{L}}\in\mathcal{P}$ with $p_{\mathcal{D}}\neq p_{\mathcal{L}}$ and $p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. In other words, the training data distribution may differ from the data source distribution, but the map connecting a data location ${\bm{x}}\in\mathcal{X}$ and its associated label value is independent on how the data is selected. Furthermore, we study a typical scenario involving molecular prediction tasks in which the data source distribution $p_{\mathcal{D}}$ is unknown, and we have only access to the locations $\mathcal{D}_{\mathcal{X}}:=\{{\bm{x}}_{i}\}_{i=1}^{n}$ of the dataset $\mathcal{D}$. Moreover, we are interested in applications in which the labeling process is computationally expensive, therefore, given a budget $b\in\mathbb{N}$ of points to label, the goal is to select a subset $\mathcal{L}_{\mathcal{X}}\subset\mathcal{D}_{\mathcal{X}}$ and label it to obtain a subset $\mathcal{L}\subset\mathcal{D}$ that solves the following optimization problem:
$\min_{\begin{subarray}{c}\mathcal{L}\subset\mathcal{D},\\ |\mathcal{L}|=b\end{subarray}}\mathbb{E}_{p_{\mathcal{D}}}[l({\bm{X}},Y,m_{\mathcal{L}})].$ (1)
@@
The optimization problem addressed in Climaco and Garcke (2024) and the one we define in (1) are similar in the sense that both focus on minimizing an expected value of the error function by selecting a subset $\mathcal{L}$ from a set of available data points $\mathcal{D}$. The key distinction lies in their objectives: in  Climaco and Garcke (2024), the goal is to minimize the maximum expected error over a finite set of known data locations. Differently, in (1), the objective is to minimize the average error, defined as the expected value of the error function over the data source distribution $p_{\mathcal{D}}$ on $\mathcal{X}\times\mathcal{Y}$.